simplify
(9^3/2) ^-2

Answer by jim_thompson5910(35256)


[9^(3/2)]^(-2) = 9^[(3/2)*(-2)]
[9^(3/2)]^(-2) = 9^(-3)
[9^(3/2)]^(-2) = 1/(9^3)
[9^(3/2)]^(-2) = 1/729